# Detecting Invariant Manifolds in ReLU-Based RNNs

### v1

@@ 5 Homo-/heteroclinic orbits and detection of chaos
The stable and unstable manifolds can be used to identify homoclinic and heteroclinic orbits, as defined in Sect. 3.2 (Def. 3, 4). Intersections between the stable and unstable manifolds of a saddle $\bm{p}$} lead to homoclinic points, or to heteroclinic points of two saddles ${\bm{p}}\neq{\bm{q}}$. The existence of such points inevitably gives rise to a complex fractal geometry (a so-scalled horseshoe structure, see Def. 5) and thus chaos (Wiggins, 1988). Finding such intersections is therefore highly illuminating for determining the dynamical behavior of a system. This is illustrated for a simple 22d PL map in Fig. 5A, where homoclinic intersections were identified by algorithm 1. For this 22d case, we can in fact analytically determine the presence of homoclinic points, as worked out in Appx. H.3, the results of which agree with algorithm 1. Fig. 5B illustrates the resulting chaotic attractor (which lies on the unstable manifold), Fig. 5C the bifurcation diagram as one varies the model’s bias term as a control parameter, and Fig. 5D the system’s two Lyapunov exponents across the chaotic range of $h_{1}$ (confirming the presence of ‘robust chaos’, as the Lyapunov exponents do not change across the chaotic regime (Banerjee et al., 1998)).
Figure 5: A) Stable (black) and unstable (red) manifolds of a saddle point (green dot) as identified by our algorithm. Note the homoclinic intersections between these two manifolds directly form the chaotic attractor. B) Structure of the chaotic attractor caused by these homoclinic intersections. C) Bifurcation diagram as a function of bias parameter $h_{1}$. D) Lyapunov exponents across the $h_{1}$-range for which the chaotic attractor exists.
@@ 6 Conclusions
Here we presented a novel semi-analytical algorithm for determining stable and unstable manifolds of fixed and cyclic points in ReLU-based RNNs. The algorithm analytically determines the local un-/stable manifolds within the subregion in which the cyclic point resides, and then by forward- or backward-iterating the RNN map $F_{\mathbf{\theta}}$ collects a few support points for spanning their extensions into neighboring linear subregions. These manifolds are profoundly important for studying an RNN’s dynamical repertoire, illuminating dynamical mechanisms in an underlying system reconstructed by the RNN, or the mechanisms by which an RNN solves a given task. Stable manifolds of saddle points segregate the state space into different basins of attraction, giving rise to the computationally important property of multistability (Durstewitz et al., 2000; Feudel et al., 2018). Intersections of stable and unstable manifolds, in turn, lead to homo- or heteroclinic points which produce a fractal geometry and chaos (Wiggins, 1988; Patra and Banerjee, 2018).
@@
Limitations arise in the context of chaotic dynamics, where invariant manifolds may fold into fractal structures. While points from these manifolds may still be sampled, the analytic construction through curved/planar segments spanned by support vectors breaks down, as the intricate, self-similar geometry of fractals cannot be captured this way. Nevertheless, we still may be able to retrieve some important dynamical characteristics by determining homo- or heteroclinic intersections, as discussed in Section˜5. Another limitation is that in the worst case scenario the algorithm may scale as $2^{P}$ with the number of linear subregions $P$. However, as shown in (Brenner et al., 2024a), the number of subregions utilized by trained PLRNNs quickly saturates, suggesting an at most polynomial scaling if one restricts attention to the domain explored by the data (see also Fig.19). In any case, we emphasize that this is the first algorithm for detecting un-/stable manifolds in ReLU-based RNNs.
7 Acknowledgements
This work was primarily supported by Samsung Advanced Institute of Technology, Samsung Electronics Co., Ltd. Additional funding was provided through the German Research Foundation (DFG) within the FOR-5159 research cluster (Du 354/14-1) and the collaborative research center TRR 265, subproject A06. Z.M. is also grateful to the Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR, Federal Ministry of Research, Technology and Space) for funding through project OIDLITDSM, No. 01IS24061.
@@ 7 Acknowledgements
Appendix A The PLRNN as 2d piecewise-linear map
The PLRNN is a piecewise linear map. As a simple example let us define a 2d map $F$ on $\mathbb{R}^{2}$ with one boundary, for which we have simple ground truth systems for, by:
where $X=\begin{pmatrix}x\\ y\end{pmatrix}$ is the position vector. The transformation matrices and vector are specified as follows:
Appendix B General 2d PL maps
A 2-unit PLRNN is a 2d PL dynamical system whose phase space is generally split into $4$ sub-regions by $4$ borders; see Appx. F for more details. For our exposition it suffices to focus on just one border, however. Hence studying general 2d PL maps, we can investigate dynamics of 2d PLRNNs locally near one border. Also, as shown in Appx. F, in some cases, there are only two different sub-regions divided by one border for 2d PLRNNs. This means in such situations, a 2d PLRNN exactly has the form of a 2d PL map. Therefore, here we examine dynamics of a 2d PL map.
A 2d PL map $T$ is a continuous map on $\mathbb{R}^{2}$ which is affine on each side of the line $\Sigma:=\big\{(x,y)^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{2}:x=0\big\}$:
$\begin{pmatrix}x^{(k+1)}\\[4.30554pt] y^{(k+1)}\end{pmatrix}\,=\,T(x^{(k)},y^{(k)})$
$\,=\begin{cases}T_{\mathcal{L}}(x^{(k)},y^{(k)})=\underbrace{\begin{pmatrix}a_{l}&&c\\[4.30554pt] b_{l}&&d\end{pmatrix}}_{{\bm{A}}_{\mathcal{L}}}\begin{pmatrix}x^{(k)}\\[4.30554pt] y^{(k)}\end{pmatrix}+\begin{pmatrix}h_{1}\\[4.30554pt] h_{2}\end{pmatrix};\hskip 5.69046ptx^{(k)}\leq 0\\[30.1388pt] T_{\mathcal{R}}(x^{(k)},y^{(k)})=\underbrace{\begin{pmatrix}a_{r}&&c\\[4.30554pt] b_{r}&&d\end{pmatrix}}_{{\bm{A}}_{\mathcal{R}}}\begin{pmatrix}x^{(k)}\\[4.30554pt] y^{(k)}\end{pmatrix}+\begin{pmatrix}h_{1}\\[4.30554pt] h_{2}\end{pmatrix};\hskip 5.69046ptx^{(k)}\geq 0\end{cases}$ (8)
where $a_{l},a_{r},b_{l},b_{r},c,d,h_{1},h_{2}\in\mathbb{R}$. The phase space of the map Eq. 8 is divided into two different sub-regions $\mathcal{L}:=\big\{(x,y)^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{2}:x\leq 0\big\}$ and $\mathcal{R}:=\big\{(x,y)^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{2}:x\geq 0\big\}$ by the borderline $\Sigma$. The map Eq. 8 is continuous across the border, but its Jacobian matrix is discontinuous across $\Sigma$.
Under some conditions on the system parameters, there exists a well-defined and invertible coordinate change that transforms the general 2d PL map Eq. 8 into a 2d PL normal form map [Glendinning and Simpson, 2021, Simpson, 2022]. For instance when $T$ is generic in the sense that $T(\Sigma)$ intersects $\Sigma$ at a unique point which is not a fixed point of $T$, or equivalently $c\neq 0$ and $(1-d)\,h_{1}+c\,h_{2}\,\neq\,0$; see [Glendinning and Simpson, 2021] for more details. However, it is not always possible to apply such a transformation. Therefore, to study 2d PLRNNs in a comprehensive way, we need to investigate the map Eq. 8 in a general case, without any restrictions on the parameters, where here we focus particularly on the case c=0c=0\, and/or $(1-d)\,h_{1}+c\,h_{2}\,=\,0$.
Appendix C Stable manifolds with complex or degenerate eigenvalues
2d spiral
For complex-conjugate eigenvalues in 2d, the associated eigenvectors span a two-dimensional invariant plane in which the motion forms a spiral (or rotation with expansion/ contraction) around the equilibrium. That is, within each linear subregion, the solution on an un-/stable manifold evolves in a straight or spiraling fashion. The spiral structure can be derived from the eigenvalues as follows. For simplicity we shift the coordinate system such that $h=0$. Suppose ${\bm{A}}+{\bm{W}}{\bm{D}}_{t}$ has a complex eigenvalue $\lambda=r\,e^{i\theta}\quad(\text{where }|\lambda|=r<1)$ with corresponding eigenvector ${\bm{v}}={\bm{A}}\;+\;i\,\bm{b},\quad{\bm{A}},\bm{b}\in\mathbb{R}^{d}.$ A solution in the direction of this eigenvector evolves as
${\bm{z}}_{t}\;=\;({\bm{A}}+{\bm{W}}{\bm{D}}_{t})^{t}\,{\bm{z}}_{0}.$
If ${\bm{z}}_{0}$ lies in the span of ${\bm{A}},\bm{b}$, then
${\bm{z}}_{t}\;=\;\mathrm{Re}\,\bigl(\lambda^{t}\,v\bigr)\;=\;r^{t}\Bigl[\cos\bigl(t\,\theta\bigr.)\,{\bm{A}}\;-\;\sin\bigl(t\,\theta\bigr.)\,\bm{b}\Bigr].$ (9)
General case
Consider a linear system of the form ${\bm{z}}_{t+1}={\bm{L}}{\bm{z}}_{t}$.
@@ Appendix C Stable manifolds with complex or degenerate eigenvalues
Case (2): If the geometric multiplicity of $\lambda_{i}$ is less than the algebraic multiplicity, then the eigenvalue is said to be defective and $\bm{L}$} admits a basis of generalized eigenvectors. In this case, $\bm{L}$} has a Jordan block ${\bm{J}}_{m}(\lambda_{i})$ of size $m>1$
${\bm{J}}_{m}(\lambda_{i})\,=\,\begin{pmatrix}\lambda_{i}&1&0&\dots&0\\ 0&\lambda_{i}&1&\dots&0\\ 0&0&\lambda_{i}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&1\\ 0&0&0&\dots&\lambda_{i}\end{pmatrix}$,
@@
${\bm{J}}^{t}_{m}(\lambda_{i})$ $\,=\,\begin{pmatrix}\lambda_{i}^{t}&&{t\choose 1}\,\lambda_{i}^{{t}-1}&&{t\choose 2}\,\lambda_{i}^{{t}-2}&&\cdots&&{t\choose m-1}\,\lambda_{i}^{{t}-m+1}\\[4.30554pt] 0&&\lambda_{i}^{t}&&{t\choose 1}\,\lambda_{i}^{{t}-1}&&\cdots&&{t\choose m-2}\,\lambda_{i}^{{t}-m+2}\\[4.30554pt] \vdots&&\vdots&&\ddots&&\ddots&&\vdots\\[4.30554pt] 0&&0&&\ldots&&\lambda_{i}^{t}&&{t\choose 1}\,\lambda_{i}^{{t}-1}\\[4.30554pt] 0&&0&&\ldots&&0&&\lambda_{i}^{t}\end{pmatrix}$.
Moreover, a generalized eigenvector ${\bm{w}}_{m}\neq 0$ of degree $m$, corresponding to the defective eigenvalue $\lambda_{i}$, satisfies
$({\bm{L}}-\lambda{\bm{I}})^{m}{\bm{w}}_{m}=0,\hskip 17.07182pt\text{but}\,\,\,({\bm{L}}-\lambda{\bm{I}})^{m-1}{\bm{w}}_{m}\neq 0,$ (11)
and $\bm{L}$} has $m$ linearly independent generalized eigenvectors associated with $\lambda_{i}$. In fact, we can construct a chain of generalized eigenvectors $\{{\bm{w}}_{1},\cdots,{\bm{w}}_{m}\}$ such that
$({\bm{L}}-\lambda{\bm{I}}){\bm{w}}_{m}={\bm{w}}_{m-1},\,\,({\bm{L}}-\lambda{\bm{I}}){\bm{w}}_{m-1}={\bm{w}}_{m-2},\,\,\cdots,\,\,({\bm{L}}-\lambda{\bm{I}}){\bm{w}}_{2}={\bm{w}}_{1},$ (12)
where ${\bm{w}}_{1}$ is a regular eigenvector. Given a chain of length $m$, the orbit contribution corresponding to the Jordan block ${\bm{J}}_{m}(\lambda_{i})$ is given by
${\bm{z}}_{t}^{defective}\,=\,\lambda_{i}^{t}\Big(c_{1}{\bm{w}}_{1}+c_{2}\,t{\bm{w}}_{2}+c_{3}\,\frac{t^{2}}{2!}{\bm{w}}_{3}+\dots+c_{m}\,\frac{t^{m-1}}{(m-1)!}{\bm{w}}_{m}\Big).$ (13)
The full orbit of the system is a linear combination of contributions from the non-defective eigenvalues $\lambda_{j},\,1\leq j\neq i\leq n$, and the defective eigenvalue $\lambda_{i}$ as
${\bm{z}}_{t}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}c_{j}\lambda_{j}^{t}{\bm{v}}_{j}+\lambda_{i}^{t}\sum_{r=1}^{m}c_{r}\frac{t^{r-1}}{(r-1)!}{\bm{w}}_{r}.$ (14)
@@
For an affine system of the form ${\bm{z}}_{t+1}={\bm{L}}{\bm{z}}_{t}+{\bm{h}}$, Eq. 14 consists of a homogeneous part (determined by the eigenstructure of $\bm{L}$}) and a particular part due to the constant bias term $\bm{h}$}. In the presence of a defective eigenvalue $\lambda_{i}$ with Jordan block of size $m$, the full orbit is given by
${\bm{z}}_{t}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}c_{j}\lambda_{j}^{t}{\bm{v}}_{j}\;+\;\lambda_{i}^{t}\sum_{r=1}^{m}c_{r}\frac{t^{r-1}}{(r-1)!}{\bm{w}}_{r}\;+\;\sum_{k=0}^{t-1}{\bm{L}}^{k}{\bm{h}}.$ (15)

### v3

@@ 5 Homo-/heteroclinic orbits and detection of chaos
The stable and unstable manifolds can be used to identify homoclinic and heteroclinic orbits, as defined in Sect. 3.2 (Def. 3, 4). Intersections between the stable and unstable manifolds of a saddle $\bm{p}$} lead to homoclinic points, or to heteroclinic points of two saddles ${\bm{p}}\neq{\bm{q}}$. The existence of such points inevitably gives rise to a complex fractal geometry (a so-scalled horseshoe structure, see Appx. I) and thus chaos due to the Smale-Birkhoff Homoclinic Theorem (wig; simpson2016unfolding). Finding such intersections is therefore highly illuminating for determining the dynamical behavior of a system. This is illustrated for a simple 22d PL map in Fig. 5A, which shows the homoclinic intersections identified by Algorithm 1. For this 22d case, we can in fact analytically determine the presence of homoclinic points, as worked out in Appx. H.2, the results of which agree with Algorithm 1. Fig. 5B illustrates the whole resulting chaotic attractor, which happens to lie on the unstable manifold in this case. Fig. 5C provides the bifurcation diagram as one varies the model’s bias term as a control parameter, and Fig. 5D the system’s two Lyapunov exponents across the chaotic range of $h_{1}$ (confirming the presence of ‘robust chaos’, as the Lyapunov exponents do not change across the chaotic regime (ben)). Finally, in Appx. Fig. 10 we show the identification of a homoclinic intersection by our algorithm in a high-dimensional empirical example, human electrocardiograms (ECG), thus confirming the presence of chaos in this signal.
Figure 5: A) Stable (black) and unstable (red) manifolds of a saddle point (green dot) and their homoclinic intersections as identified by our algorithm. B) Structure of the chaotic attractor caused by these homoclinic intersections. C) Bifurcation diagram as a function of bias parameter $h_{1}$. D) Lyapunov exponents across the $h_{1}$-range for which the chaotic attractor exists.
@@ 6 Conclusions
Here we presented a novel semi-analytical algorithm for determining stable and unstable manifolds of fixed and cyclic points in ReLU-based RNNs. Within each linear subregion, the algorithm provides an analytical construction for the manifold segments, and then recursively assembles these segments into the global un-/stable manifolds by forward- or backward-iterating the RNN map $F_{\mathbf{\theta}}$ across subregions. These manifolds are profoundly important for studying an RNN’s dynamical repertoire, illuminating dynamical mechanisms in an underlying system reconstructed by the RNN, or the mechanisms by which an RNN solves a given task. Stable manifolds of saddle points segregate the state space into different basins of attraction, giving rise to the computationally important property of multistability (dur0; feu18). Intersections of stable and unstable manifolds, in turn, lead to homo- or heteroclinic points which produce a fractal geometry and chaos (wig). While here we developed these algorithms for the class of PLRNNs (brenner_tractable_2022; hess; brenner2024alrnn), extensions to other PL systems, like SLDS (linderman_bayesian_2017; linderman_recurrent_2016) or TLS (parmelee2022core; morrison2016diversity; hahnloser2000permitted) are conceivable. SLDS, unlike PLRNNs, are not continuous across boundaries of linear regions, however, rendering them less suitable for DSR in general and making it more tricky to impose invertibility. TLNs, on the other hand, are defined in continuous rather than discrete time. Techniques for converting discrete into continuous time PLRNNs may thus be of help here (mon2). Sometimes restrictions are imposed on the weight matrices (like binary weights or mutual inhibition), leading to nice theoretical results (curto2023graph; parmelee2022core; morrison2016diversity) but also curtailing the model’s flexibility for approximating arbitrary DS.
@@
Limitations arise in the context of chaotic dynamics, where invariant manifolds may fold into fractal structures. While points from these manifolds may still be sampled, the analytic construction through curved/planar segments spanned by support vectors breaks down, as the intricate, self-similar geometry of fractals cannot be captured this way. Nevertheless, we still may be able to retrieve some important dynamical characteristics by determining homo- or heteroclinic intersections, as discussed in Section˜5. Another limitation is that in the worst case scenario the algorithm may scale as $2^{P}$ with the number of linear subregions $P$. However, as shown in (brenner2024alrnn), the number of subregions utilized by trained PLRNNs quickly saturates, suggesting an at most polynomial scaling if one restricts attention to the domain explored by the data (see also Fig. 9). Finally, although not strictly part of the present algorithm, whether one can recover all relevant manifolds also depends one whether all saddle points were detected in the first place, of course. In any case, we emphasize that this is the first algorithm for detecting un-/stable manifolds in ReLU-based RNNs.
@@
Appendix A LLM usage and Code availability
Code for the algorithms developed here and results produced is available at [upon publication]. Large Language Models (LLMs) were used for implementation of standard routines and assistance with literature search.
Appendix B Derivation of Eq. 4
The stable and unstable manifolds of saddles are invariant sets with all trajectories with initial conditions on the manifold confined to the manifold for all $t$ [alligood1996, per], $F_{\mathbf{\theta}}(W^{\sigma})\subseteq W^{\sigma}$. We recursively assemble these manifolds following Eq. 3, with points sampled from $W^{\sigma}$ following Eq. 4 within each linear subregion. Here we derive this equation starting with a linear system of the form ${\bm{z}}_{t+1}={\bm{L}}{\bm{z}}_{t}$.
@@ Appendix A LLM usage and Code availability
Case (2): If the geometric multiplicity of $\lambda_{i}$ is less than the algebraic multiplicity, then the eigenvalue is said to be defective and $\bm{L}$} admits a basis of generalized eigenvectors. In this case, $\bm{L}$} has a Jordan block ${\bm{J}}_{d}(\lambda_{i})$ of size $d>1$
${\bm{J}}_{d}(\lambda_{i})\,=\,\begin{pmatrix}\lambda_{i}&1&0&\dots&0\\ 0&\lambda_{i}&1&\dots&0\\ 0&0&\lambda_{i}&\dots&0\\ \vdots&\vdots&\vdots&\ddots&1\\ 0&0&0&\dots&\lambda_{i}\end{pmatrix}$,
@@
${\bm{J}}^{t}_{m}(\lambda_{i})$ $\,=\,\begin{pmatrix}\lambda_{i}^{t}&&{t\choose 1}\,\lambda_{i}^{{t}-1}&&{t\choose 2}\,\lambda_{i}^{{t}-2}&&\cdots&&{t\choose d-1}\,\lambda_{i}^{{t}-d+1}\\[4.30554pt] 0&&\lambda_{i}^{t}&&{t\choose 1}\,\lambda_{i}^{{t}-1}&&\cdots&&{t\choose d-2}\,\lambda_{i}^{{t}-d+2}\\[4.30554pt] \vdots&&\vdots&&\ddots&&\ddots&&\vdots\\[4.30554pt] 0&&0&&\ldots&&\lambda_{i}^{t}&&{t\choose 1}\,\lambda_{i}^{{t}-1}\\[4.30554pt] 0&&0&&\ldots&&0&&\lambda_{i}^{t}\end{pmatrix}$.
Moreover, a generalized eigenvector ${\bm{w}}_{d}\neq 0$ of degree $d$, corresponding to the defective eigenvalue $\lambda_{i}$, satisfies
$({\bm{L}}-\lambda{\bm{I}})^{d}{\bm{w}}_{d}=0,\hskip 17.07182pt\text{but}\,\,\,({\bm{L}}-\lambda{\bm{I}})^{d-1}{\bm{w}}_{d}\neq 0,$ (8)
and $\bm{L}$} has $d$ linearly independent generalized eigenvectors associated with $\lambda_{i}$. In fact, we can construct a chain of generalized eigenvectors $\{{\bm{w}}_{1},\cdots,{\bm{w}}_{d}\}$ such that
$({\bm{L}}-\lambda{\bm{I}}){\bm{w}}_{d}={\bm{w}}_{d-1},\,\,({\bm{L}}-\lambda{\bm{I}}){\bm{w}}_{d-1}={\bm{w}}_{d-2},\,\,\cdots,\,\,({\bm{L}}-\lambda{\bm{I}}){\bm{w}}_{2}={\bm{w}}_{1},$ (9)
where ${\bm{w}}_{1}$ is a regular eigenvector. Given a chain of length $d$, the orbit contribution corresponding to the Jordan block ${\bm{J}}_{d}(\lambda_{i})$ is given by
${\bm{z}}_{t}^{defective}\,=\,\lambda_{i}^{t}\Big(c_{1}{\bm{w}}_{1}+c_{2}\,t{\bm{w}}_{2}+c_{3}\,\frac{t^{2}}{2!}{\bm{w}}_{3}+\dots+c_{d}\,\frac{t^{d-1}}{(d-1)!}{\bm{w}}_{d}\Big).$ (10)
The full orbit of the system is a linear combination of contributions from the non-defective eigenvalues $\lambda_{j},\,1\leq j\neq i\leq n$, and the defective eigenvalue $\lambda_{i}$ according to
${\bm{z}}_{t}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}c_{j}\lambda_{j}^{t}{\bm{v}}_{j}+\lambda_{i}^{t}\sum_{r=1}^{d}c_{r}\frac{t^{r-1}}{(r-1)!}{\bm{w}}_{r}.$ (11)
@@
A PLRNN is not strictly linear, but affine in each subregion, see Eq. 2. For an affine system of the form ${\bm{z}}_{t+1}={\bm{L}}{\bm{z}}_{t}+{\bm{h}}$, Eq. 11 consists of a homogeneous part (determined by the eigenstructure of $\bm{L}$}) and a particular part due to the constant bias term $\bm{h}$}. In the presence of a defective eigenvalue $\lambda_{i}$ with Jordan block of size $d$, the full orbit is given by
${\bm{z}}_{t}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}c_{j}\lambda_{j}^{t}{\bm{v}}_{j}\;+\;\lambda_{i}^{t}\sum_{r=1}^{d}c_{r}\frac{t^{r-1}}{(r-1)!}{\bm{w}}_{r}\;+\;\sum_{k=0}^{t-1}{\bm{L}}^{k}{\bm{h}}.$ (12)